# On the optimality of coin-betting for mean estimation

### v2

@@ Abstract
Confidence sequences are sequences of confidence sets that adapt to incoming data while maintaining validity. Recent advances have introduced an algorithmic formulation for constructing some of the tightest confidence sequences for the mean of bounded real random variables. These approaches use a coin-betting framework, where a player sequentially bets on differences between potential mean values and observed data. This work discusses the optimality of such coin-betting formulation among algorithmic frameworks building on e-variables methods to test and estimate the mean of bounded random variables.
@@ 1 Introduction
Estimating the mean of a random variable from empirical observations is a classical problem in statistics. To account for uncertainty, a widely used approach consists in constructing a confidence set, known to contain the true mean with high probability, rather than relying solely on a point estimate. When the data are observed sequentially, one might want to update this set as new data-points become available. However, such procedure can compromise the validity of the statistical guarantee, if this was designed for a fixed sample size. To address this issue, Darling and Robbins (1967) introduced the concept of confidence sequence, a data-adaptive sequence of confidence sets whose intersection contains the desired mean with high probability.
@@
Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) have recently suggested algorithmic approaches that yield some of the tightest confidence sequences for the mean of a bounded real random variable. Both papers propose setting up a series of sequential coin-betting games, one per each mean candidate value $\mu$, where a player sequentially bets on the difference between $\mu$ and the upcoming observation. If $\mu$ matches the true mean, the game is fair, making substantial gains unlikely. A confidence sequence is then obtained by excluding those values $\mu$ that allowed the player to accumulate significant wealth.
@@
This coin-betting approach to mean estimation is a particular instance of a broader algorithmic framework for constructing confidence sequences through sequential hypothesis testing, which can be framed in terms of betting games where at each round the player has to select an e-variable (Shafer, 2021; Ramdas et al., 2022a, 2023). E-variables, non-negative random variables whose expectation is bounded by one under the tested hypothesis (Grünwald et al., 2024), have recently emerged as a powerful and increasingly popular tool for anytime-valid hypothesis testing. By serving as building blocks for constructing non-negative super-martingales, which can be seen as representing the wealth of a player in a betting game, e-variables naturally lend themselves to game-theoretic interpretations (Shafer and Vovk, 2019; Ramdas and Wang, 2024).
@@
The first part of this work focuses on round-wise testing-by-betting, a scenario where the player iteratively picks a single-round e-variable, whose choice can depend on the past observations. Our main goal is to illustrate that, when estimating the mean of a bounded real random variable, no instantiation of such round-wise testing-by-betting game can yield strictly better results than the coin-betting approach, in terms of tightness of the resulting confidence sequence. In a sense to be clarified later, coin-betting is optimal, as it represents the “simplest” formulation among those that cannot be strictly outperformed by any other such betting approach. One main novelty of this work is the introduction of a notion of optimality at the level of sets of e-variables. This perspective differs from much of the existing literature, which is primarily focused on the optimality of an individual e-variable or wealth process (e.g., log-optimality in Koolen and Grünwald 2022; Grünwald et al. 2024; Larsson et al. 2024, or admissibility in Ramdas et al. 2022a).
An alternative way to frame the first part of this work is as a characterisation of the family of single-round e-variables for testing the mean of a bounded random variable. In a similar spirit, since the appearance of the first version of this manuscript, two works have derived more general results for testing hypotheses defined by linear constraints (Clerico, 2024; Larsson et al., 2025), of which the present setting is a particular case. However, the approach we present here is direct and tailored to the simple setting considered, making it valuable for building intuition, while the more general results are significantly more abstract.
@@
We remark that testing via sequential betting with single-round e-variables is not the most general form of sequential testing with e-variables, which typically relies on multi-round e-variables and e-processes (Shafer, 2021; Koolen and Grünwald, 2022; Ramdas et al., 2022b; Ramdas and Wang, 2024). These allow for testing hypotheses over the entire data sequence, including assumptions about the dependence structure (e.g., i.i.d. or fixed conditional mean). In such cases, restricting the player to select a single-round e-variable at each step can be limiting (Koolen and Grünwald, 2022; Ramdas et al., 2022b). The second part of this work considers this broader setting. For technical convenience, we only focus on variables with finite discrete support, though we conjecture our conclusions hold more generally. We establish that when the sequence has fixed conditional mean, coin betting remains optimal even among testing methods based on multi-round e-variables and e-processes. However, we also show that this optimality result no longer holds under the more restrictive hypothesis of i.i.d. observations.
@@
We endow any Borel set $\mathcal{Z}\subseteq\mathbb{R}^{d}$ with the standard topology, and we denote as $\mathcal{P}_{\mathcal{Z}}$ the set of Borel probability measures on $\mathcal{Z}$. For $z\in\mathcal{Z}$, $\delta_{z}$ denotes the Dirac unit mass on $z$. For $P\in\mathcal{P}_{\mathcal{Z}}$ and a Borel measurable function $f$ on $\mathcal{Z}$, $\mathbb{E}_{P}[f(Z)]$ (or more compactly $\mathbb{E}_{P}[f]$) denotes the expectation of $f$ under $Z\sim P$. Given a sub-sigma-field $\mathcal{G}$, $\mathbb{E}_{P}[Z|\mathcal{G}]$ denotes the conditional expectation. Typically, $\mathcal{G}$ will be the sigma-field generated by some random variable $X$. In such case, we will write $\mathbb{E}_{P}[Z|X]$.

### v4

@@ Abstract
We consider the problem of testing the mean of a bounded real random variable. We introduce a notion of optimal classes for e-variables and e-processes, and establish the optimality of the coin-betting formulation among e-variable-based algorithmic frameworks for testing and estimating the (conditional) mean. As a consequence, we provide a direct and explicit characterisation of all valid e-variables and e-processes for this testing problem. In the language of classical statistical decision theory, we fully describe the set of all admissible e-variables and e-processes, and identify the corresponding minimal complete class.
keywords:
sequential hypothesis testing , e-variables , e-processes , mean estimation , admissibility.
journal: International Journal of Approximate Reasoning
@@ 1 Introduction
Estimating the mean of a random variable from empirical observations is a classical problem in statistics. To account for uncertainty, a widely used approach consists in constructing a confidence set, known to contain the true mean with high probability, rather than relying solely on a point estimate. When the data are observed sequentially, one might want to update this set as new data-points become available. However, such procedure may compromise the validity of the statistical guarantee, if this was designed for a fixed sample size. To address this issue, Darling and Robbins (1967) introduced the concept of confidence sequence, a data-adaptive sequence of confidence sets whose intersection contains the desired mean with high probability.
@@
Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) have recently explored algorithmic approaches that yield some of the tightest confidence sequences for the mean of a bounded real random variable. Both papers propose setting up a series of sequential coin-betting games, one per each mean candidate value $\mu$, where a player sequentially bets on the difference between $\mu$ and the upcoming observation. If $\mu$ matches the true mean, the game is fair, and substantial gains unlikely. A confidence sequence is obtained by excluding those values $\mu$ that allowed the player to accumulate significant wealth.
@@
This coin-betting approach to mean estimation is a particular instance of a broader algorithmic framework for constructing confidence sequences through sequential hypothesis testing, which can be framed in terms of betting games where at each round the player has to select an e-variable (Shafer, 2021; Ramdas et al., 2022a, 2023). E-variables, non-negative random variables whose expectation is bounded by 111 under the tested hypothesis (Grünwald et al., 2024), have recently emerged as a powerful and increasingly popular tool for anytime-valid hypothesis testing. By serving as building blocks for constructing non-negative super-martingales, which can be seen as representing the wealth of a player in a betting game, e-variables naturally lend themselves to game-theoretic interpretations (Shafer and Vovk, 2019; Ramdas and Wang, 2024).
@@
The main goal of this work is to illustrate and formalise that, when sequentially testing and estimating the (conditional) mean of a bounded real random variable, no e-variable procedure yields strictly better guarantees than the coin-betting approach. In a sense to be clarified later, coin-betting is optimal, as it represents the “simplest” formulation among those that cannot be strictly performed by any other such testing-by-betting approach. One main novelty of this work is the introduction of a notion of optimality at the level of sets of e-variables. This perspective differs from much of the existing literature, primarily focused on the optimality of an individual e-variable, or wealth process, in the betting game (e.g., log-optimality in Koolen and Grünwald 2022; Grünwald et al. 2024; Larsson et al. 2024, or admissibility in Ramdas et al. 2022a). We remark that the perspective adopted in this work can be seen as an adaptation, to the setting of e-variables, of the classical statistical problem of identifying a minimal complete class of tests (see Lehmann and Romano, 2022). Further discussion of this connection is deferred to Section 7.
@@
The first part of this work focuses on round-wise testing-by-betting, a scenario where the player iteratively picks a single-round e-variable, whose choice may depend on the past observations. This procedure naturally applies to the setting where the observations are known to be independently drawn from a fixed probability distribution, whose mean has to be estimated. However, we remark that testing via sequential betting with single-round e-variables is not the most general form of sequential testing with e-variables, which typically relies on multi-round e-variables and e-processes (Shafer, 2021; Koolen and Grünwald, 2022; Ramdas et al., 2022b; Ramdas and Wang, 2024). These tools allow for testing hypotheses over the entire data sequence, including assumptions about the dependence structure (e.g., i.i.d. or fixed conditional mean). In such cases, restricting the player to select a single-round e-variable at each step may be highly limiting (Koolen and Grünwald, 2022; Ramdas et al., 2022b). The second part of this work considers this broader setting. We establish that when the sequence has fixed conditional mean, coin-betting remains optimal even among testing methods based on multi-round e-variables and e-processes. However, we also show that this optimality result no longer holds under the more restrictive hypothesis of i.i.d. observations.
It is worth noting that an alternative way to frame the main contribution of this work is as a concrete and direct characterisation of the family of e-variables and e-processes for testing the (conditional) mean of a real bounded random variable. More precisely, the e-variables and e-processes for these tests are exactly the non-negative measurable functions or processes that are majorised by a coin-betting e-variable or e-process. Since the coin-betting formulation provides a very explicit expression for these objects, our results directly yield a fully explicit description of the full set of e-variables and e-processes for the problem at hand. Following a first pre-print of this manuscript, general characterisations for the e-variables when testing hypotheses defined by linear constraints were established by Clerico (2024) and Larsson et al. (2025). These results directly imply ours for single-round e-variables (Theorem 1), which is also explicitly discussed by these works as an application. However, the approach we present here is direct and tailored to the simple setting considered, making it valuable for building intuition, while the more general results are significantly more abstract. Moreover, the two aforementioned works focus exclusively on single-round e-variables, whereas our contribution provides a complete and explicit characterisation of both e-variables and e-processes in the sequential conditional mean testing problem. For further discussion on these works, see Section 7.
@@
We endow any Borel set $\mathcal{Z}\subseteq\mathbb{R}^{d}$ with the standard topology, and we denote as $\mathcal{P}_{\mathcal{Z}}$ the set of Borel probability measures on $\mathcal{Z}$. For $z\in\mathcal{Z}$, $\delta_{z}$ is the Dirac unit mass on $z$. For $P\in\mathcal{P}_{\mathcal{Z}}$ and a Borel measurable function $f$ on $\mathcal{Z}$, $\mathbb{E}_{P}[f(Z)]$ (or more compactly $\mathbb{E}_{P}[f]$) denotes the expectation of $f$ under $Z\sim P$. Given a sub-sigma-field $\mathcal{G}$, $\mathbb{E}_{P}[Z|\mathcal{G}]$ is the conditional expectation. We will be interested in the case of $\mathcal{G}$ being the sigma-field generated by some random variable $X$. In such case, we write $\mathbb{E}_{P}[Z|X]$.